OpenStudy (anonymous):

If (x-n)(nx+5)=nx^2-4x+k, what is n+k?

OpenStudy (campbell_st):

expand the left and side and then equate the coefficients \[(x-n)(nx+5) = nx^2 +(5 - n^2)x - 5n\] so equating coefficients on the right hand side \[(5 - n^2) = - 4\] and -5n = k add n to both sides -4n = k + n